# Strategic Classification with Non-Linear Classifiers

### v1

@@ Abstract
In strategic classification, the standard supervised learning setting is extended to support the notion of strategic user behavior in the form of costly feature manipulations made in response to a classifier. While standard learning supports a broad range of model classes, the study of strategic classification has, so far, been dedicated mostly to linear classifiers. This work aims to expand the horizon by exploring how strategic behavior manifests under non-linear classifiers and what this implies for learning. We take a bottom-up approach showing how non-linearity affects decision boundary points, classifier expressivity, and model classes complexity. A key finding is that universal approximators (e.g., neural nets) are no longer universal once the environment is strategic. We demonstrate empirically how this can create performance gaps even on an unrestricted model class.
@@ 1 Introduction
Strategic classification considers learning in a setting where users modify their features in response to a deployed classifier to obtain positive predictions [4, 3, 16]. This setting aims to capture a natural tension that can arise between a system aiming to maximize accuracy, and its users who seek favorable outcomes. Such a tension is well-motivated across diverse applications including loans approval, university admissions, job hiring, welfare and education programs, and medical services. The field has drawn much attention since its introduction, leading to advances along multiple learning fronts such as generalization [31, 34, 6], optimization [24], loss functions [25, 23], and extensions [14, 18, 11, 28].
@@
However, despite over a decade of efforts, research in strategic classification remains focused predominantly on learning with linear classifiers. Given the inherent challenges of learning in a strategic environment, linearity has the benefit of inducing a simple form of strategic behavior that permits tractable analysis. Because many realistic applications—including those mentioned above—make regular use of non-linear models, we believe that it is important to establish a better understanding of how non-linearity shapes strategic behavior, and how behavior in turn affects learning. Our paper therefore aims to extend the study of strategic classification to the non-linear regime in the hopes of developing the necessary understanding for future learning in this setting.
@@
Our first observation is that non-linear classifiers induce qualitatively different behavior than linear ones in the strategic setting. Consider first the structure induced by standard linear classifiers. Generally, users in strategic classification are assumed to best-respond to the classifier by modifying their features in a way that maximizes utility (from prediction) minus costs (from modification). For a linear classifier $h$ coupled with the common choice of an $\ell_{2}$ cost, modifications $x\mapsto x^{h}$ manifest as a projection of $x$ onto the linear decision boundary of $h$ if this incurs less than a unit cost. This simple form has several implications. First, note that all points ‘move’ in the same direction, making strategic behavior, in effect, uni-dimensional. Second, the region of points that move forms a band of unit distance on the negative side of $h$. This means that the set of points classified as positive forms a halfspace, which implies that the ‘effective’ classifier, given by $h_{\Delta}(x)\coloneqq h(x^{h})$, is also linear. Third, and as a result, the induced effective model class remains linear.
@@
Together, the above suggest that for linear classifiers coupled with $\ell_{2}$ costs, the transition from the standard, non-strategic setting to a strategic one requires anticipating which points move, but otherwise remains similar. One known implication is that the strategic and non-strategic VC dimension of linear classifiers are the same [31]. Furthermore, the Bayes-optimal strategic classifier can be derived by taking the optimal non-strategic classifier and simply increasing its bias term by one [27]. Note this means that any linearly separable problem remains separable, and that strategic classifiers can, in principle, attain the same level of accuracy as their non-strategic counterparts. Thus, while perhaps challenging to solve in practice, the intrinsic difficulty of strategic linear classification remains unchanged.
@@
The above, however, is unlikely to hold for non-linear classifiers: once the decision boundary is non-linear, different points will move in different directions. In addition to the possible distortion of the shapes of each individual classifier, the effective (i.e., strategic) model class may also change, which can have concrete implications on expressivity, generalization, optimization, and attainable accuracy. This motivates the need to explore how strategic responses shape outcomes in non-linear strategic classification, which we believe is crucial as a first step towards developing future methods. Our general approach is to examine—through analysis, examples, and experiments—how objects in the original non-strategic task map to corresponding objects in the induced strategic problem instance. Initially, our conjecture was that strategic behavior acts as a ‘smoothing’ operator on the original decision boundary, but as we show, the actual effect of strategic behavior turns out to be more involved.
@@
Taking a bottom-up approach, we study three mappings between (i) individual points, (ii) classifiers, and (iii) model classes. For points, we begin by showing that the mapping in not necessarily bijective. As a result, some points on the original decision boundary become ‘wiped out’, while others are ‘expanded’ to become intervals (arcs for $\ell_{2}$ costs) on the induced boundary. We demonstrate several mechanisms that can give rise to both phenomena. For classifiers, we show how the effects on individual points aggregate to modify the entire decision boundary. A key finding is that certain decision boundaries are impossible to express under strategic behavior. We give several examples of reasonable classifier types that cannot exist under strategic behavior, and show that some effective classifiers $h_{\Delta}$ can be traced back to (infinitely) many original classifiers $h$. The main takeaway is that universal approximators, such as neural networks or RBF kernels, are no longer universal in a strategic environment.
@@
For model classes, our results build on the observation that the mapping from original to effective classifiers within the class is non-bijective. In particular, because many $h$ can be mapped to the same $h_{\Delta}$, the VC dimension of the induced model class can shrink even from $\infty$ down to 0, causing model classes that were not learnable in the standard setting to become learnable in the strategic one. That said, the VC dimension of the induced class can generally either grow, or shrink. We highlight some conditions that guarantee either of these outcomes, though we conjecture that there is a bound on how much complexity can grow. We conclude with experiments that shed light on how strategic behavior may alter the effective complexity and maximum accuracy of the learned model class in practice.
@@ 2 Related work
Our work focuses on the common setting of supervised binary classification, where non-linear models are highly prevalent in general. This follows the original formulation of strategic classification [3, 16], which has since been extended also to online settings [8, 5, 1] and to regression [29, 2]. As noted, linear models have thus far been at the forefront of the field, with many works relying (explicitly or implicitly) on their properties or induced structure [5, 24, 10, 27, 18]. Even for works with results agnostic to model class, linear classifiers remains the focus of special-case analyses, examples, and empirical evaluation [19, 35, 25], and their generality does not shed light on the impact of non-linearity. Other works offer only tangential results, such as welfare analysis in the non-linear setting [33], but do not paint a picture of the general learning problem. Technically related, the field of offset curves deals with distortions of general non-linear curves, though it is ultimately more concerned with approximation and analytical properties and does not convey direct implications for learning [20, 12]. In terms of theory, both [34, 31] apply PAC analysis to strategic classification, the latter providing VC bounds for the linear class. Most relevant to us is [6] who study the general learnability of strategic learning under different feedback models. While not restricted to particular classes, their analysis depends crucially on the assumption that each user can move only to a finite set of possible points whose cardinality plays an explicit role. Our work tackles the more pervasive setting of movement governed by cost functions, and is therefore continuous. The above works focus primarily on statistical hardness; while this is also one of our aims, we are interested in establishing a broader understanding of the non-linear regime.
@@ 3 Preliminaries
We work in the original supervised setting of strategic classification [3, 16]. Let $x\in{\cal{X}}=\mathbb{R}^{d}$ denote user features, $y\in{\cal{Y}}=\{0,1\}$ their corresponding label, and assume pairs $(x,y)$ are drawn iid from some unknown joint distribution $D$. Given a training set $T=\{(x_{i},y_{i})\}_{i=1}^{m}\sim D^{m}$, learning aims for the common goal of finding a classifier $h\in H$ that maximizes expected predictive accuracy, where $H$ is some chosen model class. The unique aspect of strategic classification is that at test time, users modify their features to maximize gains in response to the learned $h$ via the best-response mapping:
$x^{h}\coloneqq\Delta_{h}(x)=\operatorname*{argmax}_{x^{\prime}}\alpha h(x)-c(x,x^{\prime})$ (1)
where $\alpha$ determines their utility from obtaining a positive prediction ${\hat{y}}=h(x)$, and $c$ is a function governing modification costs. We assume w.l.o.g. that the user gains no utility from a negative prediction. Like most works, we focus mainly on the $\ell_{2}$ cost, i.e., $c(x,x^{\prime})=\|x-x^{\prime}\|_{2}$, but discuss other costs in a subset of our results. Because no rational user will pay more than $\alpha$, $x^{h}$ will always be contained in the ball of radius $\alpha$ around $x$ denoted $B_{\alpha}(x)=\{x^{\prime}\,:\,c(x,x^{\prime})\leq\alpha\}$. It will also be useful to define the corresponding sphere, $S_{\alpha}(x)=\{x^{\prime}\,:\,c(x,x^{\prime})=\alpha\}$.
@@
Given the above, the learning objective is to maximize expected strategic accuracy, defined as:
@@
We assume w.l.o.g. that classifiers are of the form $h(x)=\mathds{1}{\{{f(x)\geq 0}\}}$ for some scalar score function $f$ and that the decision boundary is then the set of points for which $f(x)=0$. Our analysis takes a ‘bottom-up’ approach in which we first consider changes to individual points on the boundary; then ask how such local changes aggregate to affect a given classifier $h$; and finally, how these changes impact the entire model class $H$. It will therefore be useful to consider three mappings between original and corresponding strategic objects. For points, the mapping is $x\mapsto\nabla_{h}(x)=\{x^{\prime}\,:\,\Delta_{h}(x^{\prime})=x\,\wedge\,c(x,x^{\prime})=\alpha\}$, i.e. the set of points $x^{\prime}$ that, in response to $h$, move to $x$ at maximal cost. For classifiers, we define the effective classifier as $h_{\Delta}(x)\coloneqq h(\Delta_{h}(x))$ (see Fig. 1), and consider the mapping $h\mapsto h_{\Delta}$. This allows us to compare the original decision boundary of $h$ to the ‘strategic decision boundary’ of its induced $h_{\Delta}$, since both are defined w.r.t. raw (unmodified) inputs $x$. For model classes, we consider $H\mapsto H_{\Delta}\coloneqq\{h_{\Delta}\,:\,h\in H\}$, and refer to $H_{\Delta}$ as the effective model class.
@@
As noted, most works in strategic classification consider $H$ to be the class of linear classifiers, $h(x)=\mathds{1}{\{{w^{\top}x\geq b}\}}$. A known result is that $\Delta_{h}(x)$ can be expressed in closed form as a conditional projection operator: $x$ moves to $x_{\mathrm{proj}}=x-\frac{w^{\top}x-b}{\|w\|_{2}^{2}}w$ if $h(x)=0$ and $c(x,x_{\mathrm{proj}})\leq\alpha$, and otherwise remains at $x$ [9]. Note that all points that move do so in the same direction, i.e., orthogonally to $w$. Furthermore, each $x$ on the decision boundary of $h$ is mapped to the single point $\hat{x}$ at an orthogonal distance (or cost) of exactly $\alpha$. Considering this for all boundary points, we get that $h_{\Delta}$ simply shifts the decision boundary by $\alpha$, which gives $h_{\Delta}(x)=\mathds{1}{\{{w^{\top}x\geq b-\alpha}\}}$. As a result, $H_{\Delta}$ remains the set of all linear classifiers, and so $H_{\Delta}=H$.
@@
Informally stated, the main difference for non-linear classifiers is that points no longer move in the same direction (see Fig. 1). Note that while points still move to the ‘closest’ point on the decision boundary, the projection operator is no longer nicely defined. The result is that the shape of the decision boundary of $h$ may not be preserved when transitioning to $h_{\Delta}$. Moreover, and as we will show, some points on the decision boundary of $h$ will "disappear" and have no effect on the induced $h_{\Delta}$, while other points may be expanded to affect a continuum of points on $h_{\Delta}$. A final observation is that directionality is important in the strategic setting: because points always move towards the positive region, the shape of $h_{\Delta}$ for some $h$ may be very different than that of the same classifier but with "inverted" predictions, $h^{\prime}(x)=1-h(x)$. This lack of symmetry means that we cannot take for granted that the induced $H_{\Delta}$ will include both $h_{\Delta}$ and $1-h_{\Delta}$, even if $H$ does.
4 Points
To gain understanding about what determines the shape of $h_{\Delta}$ of a general $h$, we begin by considering a single point $x$ on the decision boundary of $h$ and examining how (and if) it maps to corresponding point(s) on the decision boundary of $h_{\Delta}$ through the point mapping $x\mapsto\nabla_{h}(x)$. We identify and discuss four general cases (see Fig. 2):
Figure 2: Types of point mappings. From left to right: one-to-one, direct wipeout, indirect wipeout, expansion, and collision. Black lines show original $h$, dashed grey lines indicate $h_{\Delta}$, and dotted circles show $S_{\alpha}(x)$.

### v2

@@ Abstract
In strategic classification, the standard supervised learning setting is extended to support the notion of strategic user behavior in the form of costly feature manipulations made in response to a classifier. While standard learning supports a broad range of model classes, the study of strategic classification has, so far, been dedicated mostly to linear classifiers. This work aims to expand the horizon by exploring how strategic behavior manifests under non-linear classifiers and what this implies for learning. We take a bottom-up approach showing how non-linearity affects decision boundary points, classifier expressivity, and model class complexity. Our results show how, unlike the linear case, strategic behavior may either increase or decrease effective class complexity, and that the complexity decrease may be arbitrarily large. Another key finding is that universal approximators (e.g., neural nets) are no longer universal once the environment is strategic. We demonstrate empirically how this can create performance gaps even on an unrestricted model class.
@@ 1 Introduction
Strategic classification considers learning in a setting where users modify their features in response to a deployed classifier to obtain positive predictions [4, 3, 16]. This setting aims to capture a natural tension that can arise between a system aiming to maximize accuracy and its users who seek favorable outcomes. Such a tension is well-motivated across diverse applications, including loan approval, university admission, job hiring, welfare and education programs, and medical services. The field has drawn much attention since its introduction, leading to advances along multiple learning fronts such as generalization [33, 36, 6], optimization [25], loss functions [26, 24], and extensions [14, 18, 11, 30].
@@
However, despite over a decade of effort, research in strategic classification remains predominantly focused on linear classification. Due to the inherent challenges of strategic learning, this focus has been reasonable, given that linearity has the benefit of inducing a simple form of strategic behavior that permits tractable analysis. Yet, since many realistic applications—including those mentioned above—make regular use of non-linear models, we believe that it is important to establish a better understanding of how non-linearity shapes strategic behavior, and how behavior in turn affects learning. Our paper therefore aims to extend the study of strategic classification to the non-linear regime.
@@
Our first observation is that non-linear classifiers induce qualitatively different behavior than linear ones in the strategic setting. Consider first the structure induced by standard linear classifiers. Generally, users in strategic classification are assumed to best-respond to the classifier by modifying their features in a way that maximizes utility (from prediction) minus costs (from modification). For a linear classifier $h$ coupled with the common choice of an $\ell_{2}$ cost, modifications $x\mapsto x^{h}$ manifest as a projection of $x$ onto the linear decision boundary of $h$ as long as it incurs less than a unit cost. This simple form has several implications. First, note that all points ‘move’ in the same direction, making strategic behavior, in effect, one-dimensional. Second, the region of points that move forms a band of unit distance on the negative side of $h$. This means that the set of points classified as positive forms a halfspace, which implies that the ‘effective classifier’, given by $h_{\Delta}(x)\coloneqq h(x^{h})$, is also linear. Third, and as a result, the induced effective model class remains linear.
@@
Together, the above suggest that for linear classifiers coupled with $\ell_{2}$ costs, the transition from the standard, non-strategic setting to a strategic one requires anticipating which points move, but otherwise remains similar. One known implication is that the strategic and non-strategic VC dimension of linear classifiers are the same [33]. Furthermore, the Bayes-optimal strategic classifier can be derived by taking the optimal non-strategic classifier and simply increasing its bias term by one [29]. Note that this means that any linearly separable problem remains separable, and that strategic classifiers can, in principle, attain the same level of accuracy as their non-strategic counterparts. Thus, while perhaps challenging to solve in practice, the intrinsic difficulty of strategic linear classification remains unchanged.
@@
The above, however, is unlikely to hold for non-linear classifiers: once the decision boundary is non-linear, different points will move in different directions. This lends to the possible distortion of the shape of a given classifier, which in turn can change the effective (i.e., strategic) model class. Such changes can have concrete implications on expressivity, generalization, optimization, and attainable accuracy. This motivates the need to explore how strategic responses shape outcomes in non-linear strategic classification, which we believe is crucial as a first step towards developing future methods. Our general approach is to examine—through analysis, examples, and experiments—how objects in the original non-strategic task map to corresponding objects in the induced strategic problem instance. Initially, our conjecture was that strategic behavior acts as a ‘smoothing’ operator on the original decision boundary, but as we show, the actual effect of strategic behavior turns out to be more involved.
@@
To build an intuition of the mechanisms at play, we take a bottom-up approach and study three mappings between (i) individual points, (ii) classifiers, and (iii) model classes. For points, we begin by showing that the mapping is not necessarily bijective. As a result, some points on the original decision boundary become ‘wiped out’, while others are ‘expanded’ to become intervals (arcs for $\ell_{2}$ costs) on the induced boundary. We demonstrate several mechanisms that can give rise to both phenomena. For classifiers, we show how the effects on individual points aggregate to modify the entire decision boundary. A key finding is that certain decision boundaries are impossible to express under strategic behavior. We give several examples of reasonable classifier types that cannot exist under strategic behavior, and show that some effective classifiers $h_{\Delta}$ can be traced back to (infinitely) many original classifiers $h$. The main takeaway is that universal approximators, such as neural networks or RBF kernels, are no longer universal in a strategic environment.
@@
Our main contribution lies in our analysis of model classes, where our results build on the observation that the mapping from original to effective classifiers within a class is not necessarily bijective. In particular, because many $h$ can be mapped to the same $h_{\Delta}$, the VC dimension of the induced model class can shrink even from $\infty$ down to 0, causing model classes that were not learnable in the standard setting to become learnable in the strategic one. That said, the VC dimension of the induced class can generally either grow or shrink. We highlight some conditions that guarantee either of these outcomes, and, in particular, provide an upper bound on the strategic complexity increase for the expressive family of piecewise-linear model classes (e.g., ReLU networks). Our analysis yields several practical takeaways for learning, model selection, and future method development (see Appendix C.10). We conclude with experiments that shed light on how strategic behavior may alter the effective complexity and maximum accuracy of the learned model class in practice.
@@ 2 Related work
Our work focuses on the common setting of supervised binary classification, where non-linear models are highly prevalent in general. This follows the original formulation of strategic classification [3, 16], which has since been extended also to online settings [8, 5, 1] and regression [31, 2]. As noted, linear models have thus far been at the forefront of the field, with many works relying (explicitly or implicitly) on their properties or induced structure [5, 25, 10, 29, 18]. Even for works with results that are agnostic to model class selection, linear classifiers remain the focus of special-case analyses, examples, and empirical evaluation [20, 37, 26], and their generality does not shed light on the impact of non-linearity. Other works offer only tangential results, such as welfare analysis in the non-linear setting [35], but do not paint a picture of the general learning problem. Our low-level analysis makes use of ideas from the field of offset curves from computational geometry [21, 12]. Whereas this field is mostly concerned with the approximation and analytical properties of individual offset curves, we are interested in the effects on class complexity and what this implies for learning. In terms of theory, both [36, 33] apply PAC analysis to strategic classification, the latter providing VC bounds for the linear class. Most relevant to us is [6], who study the general learnability of strategic learning under different feedback models. While not restricted to particular classes, their analysis considers graph costs, where users can move only to a finite set of possible points, and their results depend crucially on the cardinality of this set. Our work tackles the more pervasive setting of continuous cost functions, which imply continuous modifications. The above works focus primarily on statistical hardness; while this is also one of our aims, we are interested in establishing a broader understanding of the non-linear regime.
@@ 3 Preliminaries
We work in the original supervised setting of strategic classification [3, 16]. Let $x\in{\cal{X}}=\mathbb{R}^{d}$ denote user features, $y\in{\cal{Y}}=\{0,1\}$ their corresponding label, and assume pairs $(x,y)$ are drawn iid from some unknown joint distribution $D$. Given a training set $T=\{(x_{i},y_{i})\}_{i=1}^{m}\sim D^{m}$, learning aims for the common goal of finding a classifier $h\in H$ that maximizes the expected predictive accuracy, where $H$ is some chosen model class. The unique aspect of strategic classification is that at test time, users modify their features to maximize gains in response to the learned $h$ via the best-response mapping
${x^{h}}\coloneqq\Delta_{h}(x)=\operatorname*{argmax}_{x^{\prime}}\alpha h(x^{\prime})-c(x,x^{\prime})$ (1)
where $\alpha$ determines their utility from obtaining a positive prediction ${\hat{y}}=h(x)$, and $c$ is a function governing modification costs. We assume w.l.o.g. that the user gains no utility from a negative prediction. Like most works, we focus mainly on the $\ell_{2}$ cost, i.e., $c(x,x^{\prime})=\|x-x^{\prime}\|_{2}$, but discuss other costs in a subset of our results (see Appendix C.1 for discussion on how our results generalize to other cost functions). Because no rational user will pay more than $\alpha$ to obtain a positive prediction, $x^{h}$ will always be contained in the ball of radius $\alpha$ around $x$, denoted $B_{\alpha}(x)=\{x^{\prime}\,:\,c(x,x^{\prime})\leq\alpha\}$. It will also be useful to define the corresponding sphere, $S_{\alpha}(x)=\{x^{\prime}\,:\,c(x,x^{\prime})=\alpha\}$.
@@
Given the above, the learning objective is to maximize the expected strategic accuracy, defined as:
@@
We assume w.l.o.g. that classifiers are of the form $h(x)=\mathds{1}{\{{f(x)\geq 0}\}}$ for some scalar score function $f$, and that the decision boundary is then the surface of points for which $f(x)=0$. Our analysis takes a ‘bottom-up’ approach in which we first consider changes to individual points on the boundary, then ask how such local changes aggregate to affect a given classifier $h$, and finally, how these changes impact the entire model class $H$. It will therefore be useful to consider three mappings between original and corresponding strategic objects. For points, the mapping is ${z}\mapsto\nabla_{h}({z})=\{{x}\,:\,\Delta_{h}({x})={z}\,\wedge\,c({x},{z})=\alpha\}$, i.e., the set of points x{x} that, in response to $h$, move to z{z} at maximal cost. For classifiers, we define the effective classifier as $h_{\Delta}(x)\coloneqq h(\Delta_{h}(x))$ (see Fig. 1), and consider the mapping $h\mapsto h_{\Delta}$. Since both $h$ and $h_{\Delta}$ are defined w.r.t. raw (unmodified) inputs $x$, this allows us to compare the original decision boundary of $h$ to the ‘strategic decision boundary’ of its induced $h_{\Delta}$. For model classes, we consider the mapping $H\mapsto H_{\Delta}\coloneqq\{h_{\Delta}\,:\,h\in H\}$, and refer to $H_{\Delta}$ as the effective model class.
@@
As noted, most works in strategic classification consider $H$ to be the class of linear classifiers, $h(x)=\mathds{1}{\{{w^{\top}x\geq b}\}}$. A known result is that $\Delta_{h}(x)$ can be expressed in closed form as a conditional projection operator: $x$ moves to $x_{\mathrm{proj}}=x-\frac{w^{\top}x-b}{\|w\|_{2}^{2}}w$ if $h(x)=0$ and $c(x,x_{\mathrm{proj}})\leq\alpha$, and otherwise remains at $x$ [9]. Note that all points that move do so in the same direction, i.e., orthogonally to $w$. Furthermore, each x{x} on the decision boundary of $h$ is mapped to a single, unique point z{z} on the decision boundary of $h_{\Delta}$ at an orthogonal distance (or cost) of exactly $\alpha$. Considering this for all boundary points, we get that $h_{\Delta}$ simply shifts the decision boundary by $\alpha$, which gives $h_{\Delta}(x)=\mathds{1}{\{{w^{\top}x\geq b-\alpha}\}}$. Thus, $H_{\Delta}$ remains the set of all linear classifiers, and so $H_{\Delta}=H$.
@@
Informally stated, the main difference in the non-linear case is that points no longer move in the same direction (see Fig. 1). Note that while points still move to the ‘closest’ point on the decision boundary, the projection operator is no longer nicely defined. As a result, the shape of the decision boundary of $h$ may not be preserved when transitioning to $h_{\Delta}$. Moreover, as we will show, some points on the decision boundary of $h$ will ‘disappear’ and have no effect on the induced $h_{\Delta}$, while other points may be expanded to affect a continuum of points on $h_{\Delta}$. A final observation is that directionality is important in the strategic setting: because points always move towards the positive region, the shape of $h_{\Delta}$ for some $h$ may be very different than that of the same classifier but with "inverted" predictions, $h^{\prime}(x)=1-h(x)$. This lack of symmetry means that we cannot take for granted that the induced $H_{\Delta}$ will include both $h_{\Delta}$ and $1-h_{\Delta}$, even if $H$ does.
4 Underlying Mechanisms
Due to the added complexity of the non-linear setting, we first present some basic results and observations related to the underlying mechanisms of the non-linear setting in order to build the necessary intuition for our higher-level results. We begin by examining the point mapping ${z}\mapsto\nabla_{h}({z})$ and then turn to investigate how strategic behavior affects the classifier mapping $h\mapsto h_{\Delta}$.
4.1 Point mapping
To gain an understanding of what determines the shape of $h_{\Delta}$ of a general $h$, we begin by considering a single point z{z} on the decision boundary of $h$ and examining how (and if) it maps to corresponding point(s) on the decision boundary of $h_{\Delta}$ through the point mapping ${z}\mapsto\nabla_{h}({z})$. We identify and discuss four general cases, illustrated in Fig. 2:
Figure 2: Types of point mappings. From left to right: one-to-one (case #1), direct wipeout (case #2), indirect wipeout (case #2), expansion (case #3), and collision (case #4). Black lines show original $h$, dashed gray lines indicate $h_{\Delta}$, gold arrows show ${z}\mapsto\nabla_{h}({z})$, and dotted circles show $S_{\alpha}(x)$